OpenStudy (anonymous):

**Let A be an invertible 'n x n' matrix and B be a 'n x m' matrix.** Find A and B so that the column space of AB and the column space of B are di fferent. ((so stuck~~))

OpenStudy (maheshmeghwal9):

First of all product AB is defined or not; Let's see: - \[n \times \color{red}{\cancel n} = \color{red}{\cancel n} \times m.\]yes it is defined. but of what order would AB be: ? Let's see: - \[[n] \times \color{red}{\cancel n} = \color{red}{\cancel n} \times [m].\]It is \[[n \times m]\]As u can see.

OpenStudy (maheshmeghwal9):

So it is a question based on "rank of the matrix". I didn't noticed it before. Actually i also haven't studied the rank stuff yet so really sorry! :)

OpenStudy (maheshmeghwal9):

May be it is of some help:) http://en.wikipedia.org/wiki/Column_space